A person has $ 100 to spend on two goods X and Y whose respective prices are $3 and $5.


A. Draw the budget line.


B. What happens to the original budget line if the budget falls by 25%?


C. What happens to the original budget line if the price of X doubles?


D. What happens to the original budget line if the price of Y falls to $4?


1
Expert's answer
2022-01-28T13:10:25-0500

Solution:

A.


"p_x=3, G_x=\frac {100}{3}=33\frac{1}{3}""p_y=5, G_y=\frac{100}{5}=20"


B.



"100-0.25\times 100=75""p_x=3, G_x=\frac {75}{3}=25""p_y=5, G_y=\frac{75}{5}=15"


Its slope will remain same, but starting point, final point and all other points are lower by 25%.

C.



"p_x=6, G_x=\frac {100}{6}=16\frac{2}{3}"


Its slope is decreasing at a higher rate than previous. Starting point is same, but ending point will be different.

D.


"p_y=5, G_y=\frac{100}{4}=25"


Its slope is decreasing at a higher rate than previous. Starting point is different, but ending point will be same.